# The Additive Bachelier model with an application to the oil option market in the Covid period

### v2

@@ 5 Conclusions
  • A closed-form solution for European options pricing, based on the Lewis formula.
@@
The calibration results (discussed in Section 4) highlight the robustness of the proposed model over a very volatile period as the first period of the Covid-19 pandemic, ensuring a stable calibration for different value dates in the period from April 2020 to August 2020, and reproducing quite closely key market features.
@@ Acknowledgements
Appendix A: Bachelier formula, Lévy Bachelier and its generalization
@@
it is symmetric w.r.t the moneyness-degree.
@@
with $\hat{\eta}\in\mathbb{R}$ and $\hat{\sigma},\hat{k}\in\mathbb{R}^{+}$. This model is the generalization to the linear (Bachelier) case of the Normal Tempered Stable exponential Lévy model that can be found in excellent textbooks (see, e.g., Cont and Tankov 2003).
This model, unfortunately, is not able to reproduce adequately the observed volatility surface, as also shown in Section 4. However, a Lévy model that is calibrated on one single maturity at a time (called slice-by-slice in the text) is able to calibrate in an excellent way even the oil derivative market in the most turbulent period that has been observed so far: the first months of the Covid pandemic. Clearly, this model has a set of parameters that vary at each maturity and the model is not necessarily well posed.
For this reason, we would like to consider the natural inhomogeneous extension of the Lévy process case (21), where the parameters are time dependent, as described in the following Lemma.
The proof of the following Lemma is analogous to the one of Theorem 2.1 in Azzone and Baviera (2022, Thm.2.1), valid for exponential Additive processes.
@@
is Additive when the following conditions hold true:
@@
    $\begin{array}[]{lcl}p^{\pm}_{t}&:=&\displaystyle\frac{1}{\sigma_{t}\sqrt{t}}\left\{\displaystyle\mp\eta_{t}+\sqrt{\eta_{t}^{2}+2\,\frac{1-\alpha}{k_{t}}}\right\}\\[8.53581pt] p^{*}_{t}&:=&\displaystyle-\frac{\sigma_{t}\sqrt{t}}{k_{t}^{(1-\alpha)/\alpha}}\sqrt{\eta_{t}^{2}+2\,\frac{1-\alpha}{k_{t}}}\end{array}$
@@
The condition on $p_{t}^{\pm}$ implies that (23) is not decreasing wrt $t$ $\forall x\in\mathbb{R}$. The condition on $p_{t}^{\pm}$ implies also that
@@
We can now check whether the triplet satisfies the conditions in Sato (1999).
@@
$\sigma_{t}=\hat{\sigma},\,\eta_{t}=\hat{\eta}\sqrt{t},\,k_{t}=\frac{\hat{k}}{t}\;\;.$
In this case, condition $1$ of Lemma 5.1 is satisfied because $p^{\pm}_{t}$ are constant in time, while $p^{*}_{t}$ is decreasing, and condition $2$ holds; thus, the process is Additive. Moreover, the process is time-independent because the log-characteristic function (21) is linear in $t$.
The Additive Bachelier presented in this paper is another particular case of the Additive process (22) that, even being very parsimonious, i) presents the interesting analytical properties described in this study and ii) is able to obtain calibration results quite close to the more general (but not necessarily well posed) slice-by-slice that has time dependent parameters that vary at each maturity. In this case, $\sigma_{t}$ is s.t i) $\sigma_{t}^{2}t$ is monotone (property naturally satisfied by the ATM vol in real markets), ii) $\sigma_{t}^{2}t$ goes to zero as $t\to 0^{+}$, and
$\eta_{t}=\eta,\,k_{t}=k\;\;.$
This observation allows us to prove Proposition 2.1, as shown in Appendix B.
Appendix B: Proofs
@@
is s.t. it can be computed using iterated expectations (see, e.g., McNeil et al. 2015, Sect.3.2.2, p.77)
Remark. Thanks to Thm.3.1, p.12 of Lukacs (1972), the characteristic function of $z$ is analytic in the horizontal strip $(-p^{-},p^{+})$ with
$p^{\pm}=\mp\eta+\sqrt{\eta^{2}+2\frac{1-\alpha}{k}}\,\,.$ (28)
@@
Thus, the characteristic function of $f_{t}$ is analytic in the horizontal strip $(-p_{t}^{-},p_{t}^{+})$ with
$p_{t}^{\pm}=\frac{1}{\sigma_{t}\sqrt{t}}\,p^{\pm}\,\,.$ (29)
@@
The proof reminds the one of Lewis (2001) for exponential Lévy processes, adapted to an underlying described by (1). The call option price (6) is
$C(x,t;{\bf p})=B_{0}\,\mathbb{E}_{0}\left[f_{t}-x\right]^{+}\,\,.$
@@
Let $x\in\mathbb{R}$, the moneyness. The two parts of the Fourier transform of the call option payoff $\omega(f_{t}):=[f_{t}-x]^{+}$ are:
@@
$\mathbb{E}[f_{t}-x]^{+}=\frac{1}{2\pi}\int^{+\infty}_{-\infty}\phi_{t}(\xi)\,\frac{e^{-i\xi x}}{-\xi^{2}}\,\mathrm{d}\xi\qquad\Im(\xi)<0$
@@
Because the characteristic function $\phi_{t}(\xi)$ is analytic for $\Im(\xi)\in(-p_{t}^{-},p_{t}^{+})$, applying the Chauchy theorem, we obtain
$\mathbb{E}[f_{t}-x]^{+}=\frac{\displaystyle e^{x\,a}}{2\pi}\int^{+\infty}_{-\infty}\phi_{t}\left(\xi+{i}\,a\right)\,\frac{e^{-i\xi x}}{\left(i\,\xi-a\right)^{2}}\,\mathrm{d}\xi\,\,$,
where $a\in(-p_{t}^{-},0)$. In the case we want to consider $a\in[0,p_{t}^{+})$, we apply the residue Theorem, as in Lee (2004b), Thm.5.1 ($a>0$) and Thm.5.2 ($a=0$).
For $a\in(0,p_{t}^{+})$,
$\mathbb{E}[f_{t}-x]^{+}=-x+\frac{\displaystyle e^{x\,a}}{2\pi}\int^{+\infty}_{-\infty}\phi_{t}\left(\xi+{i}\,a\right)\,\frac{e^{-i\xi x}}{\left(i\,\xi-a\right)^{2}}\,\mathrm{d}\xi\,\,$,
and for $a=0$
$\mathbb{E}[f_{t}-x]^{+}=-\frac{x}{2}+\frac{\displaystyle e^{x\,a}}{2\pi}\int^{+\infty}_{-\infty}\left(\phi_{t}\left(\xi+{i}\,a\right)\,\frac{e^{-i\xi x}}{\left(i\,\xi-a\right)^{2}}+\frac{1}{\xi^{2}}\right)\,\mathrm{d}\xi$
@@
The first condition comes from the definition of $I(y)$ (the Bachelier formula is monotone in $y$, and the integral is well defined, so the equation for $I(y)$ is well defined).
@@
where the last equality appears in Appendix A; this proves that $I(y)\in\mathscr{C}^{1}(\mathbb{R})$.
@@
$I^{\prime}(y)=-\left.\frac{\frac{\partial\mathscr{G}(y,I)}{\partial y}}{\frac{\partial\mathscr{G}(y,I)}{\partial I}}\right|_{I=I(y)}\;\;,$
@@
Remark. It is possible to compute the derivative of the Implied vol $I(y)$ at any order $n$ via the following equation
@@
Because the value of Implied vol that solves both equations is unique, we have that
@@
that is an odd function of $\eta$, and it is equal to zero only if $\eta=0$. This proves the only if part. ∎
@@
$a\,\eta\,\sigma_{t}\sqrt{t}-\frac{a^{2}}{2}\sigma_{t}^{2}\,t>-\frac{(1-\alpha)}{k}\quad\Leftrightarrow\quad a\in(-p_{t}^{-},p_{t}^{+})\,\,$,
@@
$p_{t}^{\pm}=\frac{1}{\sigma_{t}\sqrt{t}}\left\{\mp\eta+\sqrt{{\eta}^{2}+2\,\frac{(1-\alpha)}{k}}\right\}=\frac{1}{\sigma_{t}\sqrt{t}}\,p^{\pm}\;\;.$
@@
$\lim_{x\rightarrow\pm\infty}\frac{\mathscr{I}_{t}(x)^{2}}{|x/\sqrt{t}|}=\frac{1}{2\,p_{t}^{\mp}\sqrt{t}}\,\,$,
because the characteristic function of returns $\zeta_{t}:=\frac{f_{t}}{\sqrt{t}}$ is equal to $\phi_{t}\left(\frac{u}{\sqrt{t}}\right)$ and it is analytic in the horizontal strip $(-p_{t}^{-}\sqrt{t},p_{t}^{+}\sqrt{t})$.
@@
$\lim_{y\rightarrow\pm\infty}\frac{I(y)^{2}}{|y|}=\frac{1}{\sigma_{t}}\lim_{x\rightarrow\pm\infty}\frac{\mathscr{I}_{t}(x)^{2}}{|x/\sqrt{t}|}=\frac{1}{2\,p_{t}^{\mp}\sigma_{t}\,\sqrt{t}}=\frac{1}{2\,p^{\mp}}$
@@
The proposed model is naturally sticky delta, i.e., keeping constant the model parameters $\bf p$}, the whole volatility surface as a function of the moneyness does not change ∎

### v3

@@ 5 Conclusions
  • A closed-form solution for European options pricing, the Bachelier-Lewis formula.
@@
The calibration results (discussed in Section 4) highlight the robustness of the proposed model over a very volatile period as the first period of the Covid-19 pandemic, ensuring a stable calibration for different value dates in the whole period from April 2020 to August 2020, and reproducing quite closely key market features.
While this study focuses on the WTI oil market during the Covid period, the proposed modelling framework admits natural extensions to other asset classes and market conditions, which are left as future developments for this research. First, its ability to handle negative prices makes it suitable for other commodities such as natural gas (e.g., CME 2020) and electricity (e.g., ICE Futures U.S. 2024), which can experience extreme downside volatility or near-zero prices. Second, given its roots in the Bachelier framework, the model is particularly well-suited for interest rate derivatives (e.g., swaptions, caps/floors) in low or negative interest rate environments, offering a flexible volatility smile description beyond the classic normal model. Furthermore, the parsimony of the model –relying on only a few parameters– suggests potential scalability to multi-curve or multi-commodity settings, where calibration speed is crucial. Finally, the additive process structure naturally accommodates time-inhomogeneity, paving the way for dynamic extensions that can capture regime shifts, a relevant feature for energy markets undergoing structural changes.
@@ Acknowledgements
Appendix Appendix : Bachelier formula, Lévy Bachelier and its generalization: Bachelier formula, Lévy Bachelier and its generalization
@@
it is symmetric wrt the moneyness degree.
@@
with $\hat{\eta}\in\mathbb{R}$ and $\hat{\sigma},\hat{k}\in\mathbb{R}^{+}$, and $\psi(\bullet;k,\alpha)$ is defined in (3). This model is the generalization to the linear (Bachelier) case of the normal tempered stable exponential Lévy model that can be found in excellent textbooks (see, e.g., Cont and Tankov 2003).
This model, unfortunately, is not able to reproduce the observed volatility surface adequately, as also shown in Section 4. However, a Lévy model that is calibrated on one single maturity at a time (also called slice-by-slice) is able to calibrate in an excellent way even the oil derivative market in the most turbulent period that has been observed so far: the first months of the Covid pandemic. Clearly, this model has a set of parameters that vary at each maturity and it is not necessarily well posed.
For this reason, we would like to consider the natural inhomogeneous extension of the Lévy process case (23), where the parameters are time dependent, as described in the following Lemma. The proof of the following Lemma is analogous to that of Theorem 2.1 in Azzone and Baviera (2022, Thm.2.1), valid for exponential additive processes, and it is provided in Appendix Appendix : Proofs.
@@
where the function $\psi$ is defined as in (3), is additive when the following conditions hold true:
@@
    $\begin{array}[]{lcl}p^{\pm}_{t}&:=&\displaystyle\frac{1}{\sigma_{t}\sqrt{t}}\left\{\displaystyle\pm\eta_{t}+\sqrt{\eta_{t}^{2}+2\,\frac{1-\alpha}{k_{t}}}\right\}\\[8.53581pt] p^{*}_{t}&:=&\displaystyle-\frac{\sigma_{t}\sqrt{t}}{k_{t}^{(1-\alpha)/\alpha}}\sqrt{\eta_{t}^{2}+2\,\frac{1-\alpha}{k_{t}}}\end{array}$
@@
$\sigma_{t}=\hat{\sigma},\,\eta_{t}=\hat{\eta}\sqrt{t},\,k_{t}=\frac{\hat{k}}{t}\;\;.$
In this case, condition $1$ of Lemma 5.1 is satisfied because $p^{\pm}_{t}$ are constant in time, while $p^{*}_{t}$ is decreasing, and condition $2$ holds; thus, the process is additive. Moreover, the process has stationary increments because the log-characteristic function (23) is linear in $t$ (see, e.g., Cont and Tankov 2003, Def.3.1, Sec.3.1).
The additive Bachelier presented in this paper is another particular case of the additive process (24) that, even being very parsimonious, i) presents the interesting analytical properties described in this study and ii) is able to obtain calibration results quite close to the more general (but not necessarily well posed) slice-by-slice, that has time dependent parameters that vary at each maturity. In the additive Bachelier, $\sigma_{t}$ is s.t. i) $\sigma_{t}^{2}t$ is monotone (property naturally satisfied by the ATM vol in real markets), ii) $\sigma_{t}^{2}t$ goes to zero as $t\to 0^{+}$, and
$\eta_{t}=\eta,\,k_{t}=k\;\;.$
This observation allows us to prove Proposition 2.1, as shown in Appendix Appendix : Proofs.
Appendix Appendix : Proofs: Proofs
@@
The condition on $p^{\pm}_{t}$ implies that (25) is not decreasing wrt $t$ $\forall x\in\mathbb{R}$. The condition on $p^{\pm}_{t}$ implies also that
@@
We can now check whether the triplet satisfies the conditions in Sato (1999, Thm.9.8, p.52).
@@
is s.t. it can be computed using iterated expectations (see, e.g., McNeil et al. 2015, Sec.3.2.2, p.77)
Remark. Thanks to Thm.3.1, p.12 of Lukacs (1972), the characteristic function of $z$ is analytic in the horizontal strip $(-p^{+},p^{-})$ with
$p^{\pm}=\pm\eta+\sqrt{\eta^{2}+2\frac{1-\alpha}{k}}\,\,.$ (30)
@@
Thus, the characteristic function of $f_{t}$ is analytic in the horizontal strip $(-p^{+}_{t},p^{-}_{t})$ with
$p^{\pm}_{t}=\frac{1}{\sigma_{t}\sqrt{t}}\,p^{\pm}\,\,.$ (31)
@@
The proof reminds the one of Lewis (2001) for exponential Lévy processes, adapted to an underlying described by (1).
@@
Let us consider the call option price (7) for $C(x,t;{\bf p})$. Let $x\in\mathbb{R}$, the moneyness. The two parts of the Fourier transform of the call option payoff $\omega(f_{t}):=[f_{t}-x]^{+}$ are:
@@
$\mathbb{E}_{0}[f_{t}-x]^{+}=\frac{1}{2\pi}\int^{+\infty}_{-\infty}\phi_{t}(\xi)\,\frac{e^{-i\xi x}}{-\xi^{2}}\,\mathrm{d}\xi\qquad\Im(\xi)<0$
@@
Because the characteristic function $\phi_{t}(\xi)$ is analytic for $\Im(\xi)\in(-p^{+}_{t},p^{-}_{t})$, applying the Chauchy theorem, we obtain
$\mathbb{E}_{0}[f_{t}-x]^{+}=\frac{\displaystyle e^{x\,a}}{2\pi}\int^{+\infty}_{-\infty}\phi_{t}\left(\xi+{i}\,a\right)\,\frac{e^{-i\xi x}}{\left(i\,\xi-a\right)^{2}}\,\mathrm{d}\xi\,\,$,
where $a\in(-p^{+}_{t},0)$. In the case we want to consider $a\in[0,p^{-}_{t})$, we apply the residue Theorem, as in Lee (2004a), Thm.5.1 ($a>0$) and Thm.5.2 ($a=0$, Cauchy principal value).
For $a\in(0,p^{-}_{t})$,
$\mathbb{E}_{0}[f_{t}-x]^{+}=-x+\frac{\displaystyle e^{x\,a}}{2\pi}\int^{+\infty}_{-\infty}\phi_{t}\left(\xi+{i}\,a\right)\,\frac{e^{-i\xi x}}{\left(i\,\xi-a\right)^{2}}\,\mathrm{d}\xi\,\,$,
and for $a=0$ (where the integral is intended as its principal value)
$\mathbb{E}_{0}[f_{t}-x]^{+}=-\frac{x}{2}+\frac{\displaystyle 1}{2\pi}\int^{+\infty}_{-\infty}\left(-\phi_{t}\left(\xi\right)\,\frac{e^{-i\xi x}}{\xi^{2}}+\frac{1}{\xi^{2}}\right)\,\mathrm{d}\xi$
@@
The first condition comes from the definition of $I(y)$ (that is well posed thanks to Proposition 2.5).
@@
where the last equality appears in Appendix Appendix : Bachelier formula, Lévy Bachelier and its generalization; this proves that $I(y)\in\mathscr{C}^{1}(\mathbb{R})$.
@@
$I^{\prime}(y)=-\displaystyle\left.\frac{\frac{\partial\mathscr{G}(y,I)}{\partial y}}{\frac{\partial\mathscr{G}(y,I)}{\partial I}}\right|_{I=I(y)}\;\;,$
@@
Remark. It is possible to compute the derivative of the implied vol $I(y)$ at any order $n$ (where it is well-defined) via the following equation
@@
Because the value of the implied vol that solves both equations is unique, we have that
@@
that is an odd function of $\eta$, and it is equal to zero only if $\eta=0$. This proves the only if part ∎
@@
$a\,\eta\,\sigma_{t}\sqrt{t}-\frac{a^{2}}{2}\sigma_{t}^{2}\,t>-\frac{(1-\alpha)}{k}\quad\Leftrightarrow\quad a\in(-p^{+}_{t},p^{-}_{t})\,\,$,
@@
$p^{\pm}_{t}=\frac{1}{\sigma_{t}\sqrt{t}}\left\{\pm\eta+\sqrt{{\eta}^{2}+2\,\frac{(1-\alpha)}{k}}\right\}=\frac{1}{\sigma_{t}\sqrt{t}}\,p^{\pm}\;\;.$
@@
$\lim_{x\rightarrow\pm\infty}\frac{\mathscr{I}_{t}(x)^{2}}{|x/\sqrt{t}|}=\frac{1}{2\,p_{t}^{\pm}\sqrt{t}}\,\,$,
because the characteristic function of returns $\zeta_{t}:=\frac{f_{t}}{\sqrt{t}}$ is equal to $\phi_{t}\left(\frac{u}{\sqrt{t}}\right)$ and it is analytic in the horizontal strip $(-p^{+}_{t}\sqrt{t},p^{-}_{t}\sqrt{t})$.
@@
$\lim_{y\rightarrow\pm\infty}\frac{I(y)^{2}}{|y|}=\frac{1}{\sigma_{t}}\lim_{x\rightarrow\pm\infty}\frac{\mathscr{I}_{t}(x)^{2}}{|x/\sqrt{t}|}=\frac{1}{2\,p^{\pm}_{t}\sigma_{t}\,\sqrt{t}}=\frac{1}{2\,p^{\pm}}$
@@
The proposed model is naturally sticky delta, i.e. keeping constant the model parameters $\bf p$}, the whole volatility surface as a function of the moneyness does not change ∎